OpenStudy (anonymous):

Consider the function f(x) = ax+3 over x-b. Find a and b given that y=f(x) has asymptotes with equations x = -1 and y = 2.

OpenStudy (anonymous):

so at x = -1, the equation goes to infinity. that means the denominator goes to zero. x - b = 0 -1 - b = 0 b = -1 the limit as x approaches infinity is y=2. since the "3" and the "b" become meaningless at x = inf, we're left with: ax / x = 2 a = 2 the graph of this equation is: http://bit.ly/qrWgCB

OpenStudy (anonymous):

Thanks!